OpenStudy (kropot72):

Four cards are randomly chosen (without replacement) from a deck of 52 cards. Let K be the event that two of the four cards are King and the other two are Aces. Let E be the event that one of the four cards is a King of Hearts. The conditional probability P(K | E) is equal to ?

OpenStudy (kropot72):

I have got as far as \[P(K|E)=\frac{K \cap E)}{P(E)}\] But how do I calculate K intersection E?

OpenStudy (anonymous):

the intersection of K and E means two are aces, one is the king of hearts and one is some other king

OpenStudy (anonymous):

i think there are 18 ways to get this, \[\binom{4}{2}\times 3\]

OpenStudy (anonymous):

number of ways to get 2 aces and 2 kings is \[\binom{4}{2}\times \binom{4}{2}=6\times 6=36\]

OpenStudy (anonymous):

a simpler way to think of it is off all the possibilities with two aces and two kings, half of them have the king of hearts

OpenStudy (kropot72):

Many thanks for your explanation which is clearing my mind on this one. I will try taking it from this point. Thank you again for your attention. :)